i donot know how to slove this problem:
2x-8=10
please, step by step show me how to answer because my teacher CANOT teach

Answer by queenofit(120)

Ok first let me tell you I like to think of the numbers and variables being on opposing teams. Each "team" wants to be with it's team-mates and not on the other teams side. In every equation like this, you try to get the teams to opposite sides of the equal sign with the variable (or quarterback)alone on one side equal to the other. To do this we must have a fair game, so what ever you do to one side you must do the same to the other. Now, lets work your equation.
2x-8=10
first get all your number team on one side and your variable on the other. This may take more than one step. First you rewrite the equation adding 8 to each side to move the 8 to the number team (on the 10's side)
2x-8+8=10+8
This will leave you with the following after you combine the like terms.
2x=18
But you still have a 2 guarding the quarterback. The 2 is actually being multiplied with the x, so to get it on the right team you must divide each team
by the 2 to move it to the number's team. like this
2x/2=18/2
this way the 2's cancell on the left side of the equation leaving only X (the quarterback). Then divide the 18 by 2 and you get the following
x=9
you can check this by plugging the 9 into the equation in place of the x.
2(9)-8=10
18-8=10
see how this works?
hope this helped you